When factoring the following by the reverse FOIL method, the factors of 6 must add up to what value?
I have no clue as to where to begin.
f(x) = x^2 + 5x + 6
a. 1
b. 1+6+5=12
c. 6
d. 5

Answer by nabla(475)

Consider the following:
(1)f(x)=(x+1)(x+2)
expand it:
(2)f(x)=x^2+2x+1x+2
(3)f(x)=x^2+3x+2
Similar to the problem you proposed, we could ask "the factors of 2 must add up to what value?" What are the factors of 2? They add up to what? From (2) above, we see that 2 and 1 are the coefficients of individual x terms from the expansion. Thus, to address your problem we need to find out the factors of 6 that add up to 5. It is the same process.
Your problem factors as f(x)=(x+3)(x+2). It should be evident by now that the correct answer is d.5.